OpenStudy (anonymous):

help me prove this cosA/1+sinA = secA-tanA

OpenStudy (anonymous):

lol!!! look at your expression, what is the right hand side???

OpenStudy (anonymous):

secA-tanA ???????? lol any hpnd?

OpenStudy (anonymous):

times both numerator and denominator by (1-sin) you have \[\dfrac{cos}{1+sin}*\dfrac{1-sin}{1-sin}=\dfrac{cos -sincos}{cos^2}=\dfrac{1}{cos}-\dfrac{sin}{cos}=sec-tan\] You put A after each of the sin/cos to submit hihihih

OpenStudy (anonymous):

oh thaks frnd